OpenStudy (highschoolmom2010):

find the measure of MN to the nearest tenth using the law of cosines.

OpenStudy (highschoolmom2010):



OpenStudy (anonymous):

is there some sort of formula

OpenStudy (anonymous):

because i know that if you add all the angels it has to equal 180 to be a right triangle

OpenStudy (highschoolmom2010):

Law of Cosines relates the cosine of each angle to the side lengths of the triangle. it can be either \[a^2=b^2+c^2=2bccosA\] \[b^2=a^2+c^2-2accosB\] \[c^2=a^2+b^2-2abcosC\] those are the formulas I was given...

OpenStudy (anonymous):

@thomaster

OpenStudy (raden):

MN^2 = ML^2 + LN^2 - 2(ML)(LN)cos104

OpenStudy (highschoolmom2010):

how ^^

OpenStudy (raden):

so, MN^2 = (48)^2 + (29)^2 - 2(48)(29)cos104 simplify then take the square root to get MN

OpenStudy (highschoolmom2010):

can you explain it a little more

OpenStudy (highschoolmom2010):

I don't understand how to do this

OpenStudy (anonymous):

@thomaster

OpenStudy (phi):

This is a Law of Cosines problem

OpenStudy (highschoolmom2010):

yes how the heck do I do law of cosines....@phi

OpenStudy (phi):

The Law of Cosines says if you want to find the length of side c, then you need the angle opposite that side (often called C), and the two sides 

OpenStudy (phi):

to do you your problem, you need to find side MN what is the angle opposite side MN ?

OpenStudy (highschoolmom2010):

<L

OpenStudy (phi):

= 104º which we will call C (to match the formula) we also need the other two sides,which we call a and b they give us the other 2 sides. we can call one a and the other b (which is which does not matter for the formula) so let a= 48 and b= 29 we are ready! write down the formula \[ c^2 = a^2 + b^2 - 2\ a\ b\ \cos(C) \] replace the letters a, b and C with the numbers from the problem can you do that ?

OpenStudy (highschoolmom2010):

\[104^2=48^2+29^2-2(48)(29) \cos 104\] like that??

OpenStudy (phi):

exactly like that. except the left side is not 104^2 that is lower case c, representing the length of the side opposite the angle 104º so you have \[ c^2 = 48^2+29^2−2(48)(29)\cos104 \]

OpenStudy (phi):

now do the arithmetic on the right side to find c^2 take the square root to find c which is the side MN for this problem.

OpenStudy (highschoolmom2010):

oh my mistake......what do we do next

OpenStudy (phi):

get out your calculator and figure out what the right side is

OpenStudy (highschoolmom2010):

do I do -2(48)*(29)

OpenStudy (highschoolmom2010):

i got a negative number....

OpenStudy (phi):

you do what it says \[ c^2 = 48^2+29^2−2(48)(29)\cos104 \] that means find 48*48 add that to 29*29 add that to -2 * 48* 29 * cos(104)

OpenStudy (phi):

can you do -2*48*29*cos(104) ? or you can do it using google: type into google 48^2 + 29^2 - 2*48*29*cos(104 degrees)= but I would learn how to get the same number using a calculator

OpenStudy (highschoolmom2010):

3818.51055735

OpenStudy (highschoolmom2010):

that's what popped up

OpenStudy (phi):

yes, that means c^2 = 3818.51055735 see if you can get the same number using your calculator. now to find c, take the square root of both sides sqr(c^2) is c sqr(3818.51055735) = 61.794098 c= 61.794098 find the measure of MN to the nearest tenth now round to the nearest tenth

OpenStudy (highschoolmom2010):

61.8

OpenStudy (highschoolmom2010):

ty @phi